# A Survey on Large Language Models from General Purpose to Medical Applications: Datasets, Methodologies, and Evaluations

### v1

@@ II-A1 Existing Public Datasets
Question-Answer Pairs (QA) include single-turn question-answer, multi-turn question-answer (conversations), and multiple-choice question-answer datasets. Single-turn question-answer datasets, such as Huatuo-26M [62] and ChiMed [63], primarily consist of questions and answers collected from real-life scenarios. Each question addresses a specific topic within the medical domain, enabling medical LLMs to learn a specialized manner and style of responding. However, due to its single-turn QA nature, it falls short of enhancing the ability of medical LLMs to process contextual information. Conversely, datasets in a multi-turn QA format, such as MedDialog [64], BianqueCorpus [39], and CMtMedQA [38], have the potential to augment the capacity of medical LLMs to handle contextual information and simulate authentic doctor-patient interactions. Additionally, these multi-round QA datasets often include instances where doctors inquire about patients’ conditions, thereby bolstering the medical LLMs’ ability to ask proactive questions and guide users in describing their situations. Multiple-choice QA datasets, such as MMLU [65], MedQA [66], PubMedQA [67], and MedMCQA [68], consist of questions and options, with some derived from actual medical exam questions. In addition to correct answers, some of these datasets include detailed explanations. This enhances the model’s ability to select correct answers and improves its understanding and reasoning about complex medical information.
@@
Human Preference Data, such as Zhongjing_rlhf [38] and MedicalGPT [35], consist of a question, a chosen response, and a rejected response. The chosen response aligns with human values (more like a doctor’s response), whereas the rejected response is not consistent with human values compared to chosen response. These datasets typically require manual labeling or verification, resulting in a smaller number of public datasets of this type compared to other types. Such datasets are used in the human alignment training phase of medical LLMs to ensure that the model’s responses align with human linguistic conventions and exhibit characteristics that are doctor-like, patient-friendly, harmless, and professional.
@@
Knowledge Graphs represent concepts, entities, and their interrelationships in the real world in a structured format, aligning information more closely with human cognition. They enhance the organization, management, and comprehension of vast data. Medical knowledge graphs usually contain rich structured medical knowledge with low data noise. This type of data can motivate medical LLMs to more accurately process medical terminology and causal relationships, thus improving models’ performance in reasoning medical tasks. Common medical knowledge graphs include UMLS [69], CMeKG [70], and BIOS[71]. Since these are not in QA pairs or plain text format, a carefully designed data transformation pipeline is required to convert them.
@@
NLP Task datasets include Named Entity Recognition (NER), Relation Extraction (RE), Causal Relation Extraction (CRE), Text Classification (TC), Text Pair-Semantic Similarity (TP-SS), Text Pair-Textual Entailment (TP-TE), and others. These datasets enhance model performance in various medical tasks and instruction following. Me-LLaMA [50] and Taiyi [42] utilize these conventional medical NLP datasets to improve their models’ generalization capabilities. It is important to note that these data generally cannot be used directly for training medical LLMs and must be transformed into plain text or question-answer pair format by data processing.
@@
II-A2 Public Medical Corpus
@@
Publicly accessible online medical data is a primary source for training medical LLMs. This data encompasses diverse medical information from various regions, populations, and platforms. While it offers extensive coverage and a wide range of medical knowledge, it often presents challenges related to privacy and data quality, necessitating further cleaning and processing. Here, we summarize six common data sources for medical LLMs: medical encyclopedias, medical textbooks, medical academic literature, medical websites, medical guidelines, and package inserts.
@@
Medical Encyclopedias are comprehensive reference resources, typically available online, that systematically document a wide range of medical knowledge, including diseases, symptoms, diagnostic methods, treatments, and related medical technologies. Compiled by professional medical groups, academic institutions, or authoritative publishers, their content undergoes rigorous review and regular updates to ensure accuracy and currency. The structured organization of these encyclopedias facilitates efficient data retrieval and processing. Notable medical encyclopedia websites include MedlinePlus [72], Mayo Clinic [73], WebMD [74]. Medical LLMs often use encyclopedias as a training data source to enhance their medical knowledge [45, 54]. Although the format of encyclopedia content is suitable for the pretraining of these models, further data processing and transformation are necessary for instruction fine-tuning.
@@
Medical Textbooks provide a systematic introduction to medical knowledge and skills, serving as comprehensive references for medical students, physicians, nurses, and other healthcare professionals. They cover a wide range of topics from basic sciences (e.g., anatomy, physiology, biochemistry) to clinical practice (e.g., internal medicine, surgery, pediatrics, obstetrics and gynecology). Authored by experts and peer-reviewed, these textbooks ensure the accuracy and authority of medical information. Common sources include FreeBooks4Doctors [75], Internet Archive [76], and Bookboon [77]. Similar to medical encyclopedias, medical textbooks offer extensive, specialized, and structured information. Consequently, medical LLMs [46, 51, 55, 49] utilize these textbooks as training data to incorporate professional and precise medical knowledge. However, the specialized terminology and complex linguistic structures in medical textbooks can limit their accessibility for users without a medical background.
@@
Medical Academic Literature encompasses specialized publications that research, explore, and discuss various fields of medicine. This literature includes peer-reviewed journal articles, scholarly books, conference papers, and research reports. Covering specialties such as clinical medicine, biomedical sciences, public health, pharmacology, and physiology, these publications provide up-to-date medical knowledge and evidence for researchers, clinicians, students, and allied professionals. Common sources include PubMed [78], Embase [79], and ScienceDirect [80]. Due to its specialization, accuracy, and currency, medical academic literature is frequently used to train medical LLMs [44, 50, 30], enhancing their medical expertise. However, the specialized terminology and technical content can make these models less accessible to the general public. Additionally, acquiring academic literature data can present versioning challenges.
@@
Medical Websites encompass various online platforms, such as forums, consultations, news, and educational resources. Specifically, medical consultation platforms allow patients to communicate with doctors online for medical advice, diagnoses, and treatment plans, enhancing the convenience and accessibility of healthcare services. These platforms also offer doctors a flexible way to provide care more efficiently. Common medical websites include Ding Xiang Yuan [81], Ping An Good Doctor [82], Teladoc Health [83], 98point6 [84], and Haodf [85]. For training medical LLMs (LLMs), conversation data from medical consultation platforms is particularly valuable. This data is generally more accessible to the public compared to academic literature and textbooks due to its real-world, colloquial nature. Additionally, the multi-round question-and-answer format helps improve the models’ ability to process contextual information. Several models, such as MMedLM [49], ChatDoctor [25], and GPT-Doctor [47], use this type of data to align their response styles with those of real doctors. However, data from medical consultation platforms must be collected with strict adherence to privacy regulations to ensure patient information security. The quality of this data can vary, with some information being inaccurate or incomplete. Therefore, it is essential to screen and clean the data to maintain the accuracy and reliability of the medical information used for training.
@@
Package Inserts are documents that provide comprehensive information about medications and their proper use. Intended for patients, doctors, and other healthcare professionals, they ensure safe and effective medication use. Package inserts typically include details such as the drug name and ingredients, usage and dosage, indications, contraindications, warnings and precautions, drug interactions, and side effects. Common sources for drug insert data include RxList and Drugs. These documents contain rich information that can enhance medical LLMs’ understanding of drugs and support decision-making processes. However, there are limitations. Package inserts offer standardized medication instructions and may not account for individual patient conditions or concurrent medication use, which can affect the ability of medical LLMs to provide personalized healthcare and treatment recommendations.
@@
II-A3 Professional Medical Organization Corpus
@@
Data from specialized medical institutions, such as hospitals and healthcare centers, is a crucial information source for training medical LLMs. These data includes patients’ medical histories, diagnostic information, treatment plans, and doctor-patient interactions. Here, we summarize four commonly used professional medical organization corpus sources: Clinical Notes, Electronic Health Records, Medical Prescriptions, and Doctor-Patient Conversation.
@@
Clinical Notes are records created by healthcare professionals during the consultation, diagnosis, treatment, and monitoring of patients. Typically prepared by doctors, nurses, and therapists, these notes include patient information, symptom descriptions, diagnoses, assessments, treatment plans, and progress notes. MIMIC [91] is a publicly accessible database of clinical notes, offering diverse and comprehensive data covering the entire course of a patient’s medical care. These data will enhance medical LLMs’ abilities in clinical decision-making and case understanding. For instance, models like Zhongjing [38] and Me-LLaMA [50] use clinical notes for continued pretraining to enrich their knowledge and understand real-world clinical scenarios and reasoning. Meanwhile, Clinical Camel [33] aims to enable LLMs to synthesize plausible clinical records.
@@
Electronic Health Records (EHRs) are digital versions of patients’ medical records, offering a more comprehensive overview than clinical notes [92]. EHRs include medical history, diagnoses, medication regimens, immunization dates, allergy information, radiology images, and laboratory test results. This real-time, patient-centric record system provides authorized users with immediate, secure access to patient information. Publicly accessible sources for EHRs include eICU [93], MIMIC-IV [94], and CPRD [95]. Given the close relation to medical practice and the multidimensional nature of the data, some works [38, 36] have utilized EHR data to train healthcare domain models, enhancing their usefulness and authenticity. However, different healthcare organizations may use varying EHR systems and record formats, necessitating additional preprocessing to standardize the data. Furthermore, EHRs contain personal medical information that must be anonymized to protect patient privacy.
@@
Medical Prescription is an official directive from a physician or authorized healthcare professional to a pharmacist, detailing the medication to be dispensed to a patient, including the specific drug, dosage, frequency, and administration method. These data reflect real-world prescription practices, providing valuable and authentic information to enhance models’ decision-making in healthcare. Some hospitals have integrated electronic medical prescriptions into their EHR systems, significantly facilitating scientific research on this data. Qibo[55] utilizes Traditional Chinese Medicine (TCM) prescriptions as pretraining corpus to incorporate TCM knowledge into medical LLMs. Medical Prescription data is relatively scarce and generally requires collaboration with hospitals to access. Additionally, these datasets may exhibit biases, with certain hospitals or physicians favoring specific drugs or treatments. Errors, omissions, or inaccuracies, such as misspelled drug names or incorrect dosages, can also compromise the effectiveness of model training and application.
@@
Doctor-Patient Conversation data encompasses real-life QA and interactions between doctors and patients, including symptom descriptions, diagnostic information, and patient feedback. These conversations typically involve multiple rounds of communication, with doctors actively asking questions. Such data can enhance the capabilities of medical LLMs in assisting diagnosis, providing treatment suggestions, asking relevant questions, and understanding context. HuotuoGPT [8] uses real medical conversation data in serving instruction fine-tuning to enhance long conversations for medical LLMs. However, accessing this data is challenging due to privacy concerns from hospitals and patients, and the colloquial and emotional nature of these conversations requires future processing.
@@ II-A4 Synthetic Data
Synthetic data is generated by prompting a LLM to create content based on its internal knowledge or contextual information. The common general synthetic datasets are Alpaca-52k, ShareGPT, which are in QA format. These datasets aim to align one’s own LLM with state-of-the-art models at minimal cost. In the medical domain, several LLMs utilize state-of-the-art models to generate training datasets, constructing both single-round and multi-round QA formats.
@@
For single-round QA data, ChatMed-Consult [28] takes medical questions from the internet as input to ChatGPT, and then concatenates its outputs to form an instruction fine-tuning dataset. HuatuoGPT [8] employs a self-instruct [96] method, generating a medical instruction dataset by manually creating seed instructions based on specific roles and use cases. Alpacare [40] initially creates a small set of clinician-curated seed tasks, considering four dimensions: topic, view, type, and difficulty level. Each task includes detailed information specific to these dimensions, instructions, and possibly corresponding inputs. GPT-4 then generates new instructions based on these seed tasks, filtering out highly similar ones. Finally, ChatGPT answers each generated instruction to compose the MedInstruct-52k instruction fine-tuning dataset. For multi-round QA data, HuatuoGPT employs two ChatGPTs: one simulates a doctor and the other as a patient. Real data is utilized to stimulate conversations between these two ChatGPTs, resulting in a multi-round QA dataset.
@@
The quality of the generated data depends on the effectiveness of prompts. Well-designed prompts can significantly stimulate the intrinsic knowledge of LLMs, resulting in higher-quality data. However, since this data is not derived from real-world sources, manual checking and correction are typically required to enhance its quality.
II-B Data Processing

### v2

@@ 2.1.1 Existing Public Datasets
Question-Answer Pairs (QA) include single-turn question-answer, multi-turn question-answer (conversations), and multiple-choice question-answer datasets. Single-turn question-answer datasets, such as Huatuo-26M [88] and ChiMed [143], primarily consist of questions and answers collected from real-life scenarios. Each question addresses a specific topic within the medical domain, enabling medical LLMs to learn a specialized manner and style of responding. However, due to its single-turn QA nature, it falls short of enhancing the ability of medical LLMs to process contextual information. Conversely, datasets in a multi-turn QA format, such as MedDialog [185], BianqueCorpus [23], and CMtMedQA [179], have the potential to augment the capacity of medical LLMs to handle contextual information and simulate authentic doctor-patient interactions. Additionally, these multi-turn QA datasets often include instances where doctors inquire about patients’ conditions, thereby bolstering the medical LLMs’ ability to ask proactive questions and guide users in describing their situations. Multiple-choice QA datasets, such as MMLU [60], MedQA [72], PubMedQA [73], and MedMCQA [115], consist of questions and options, with some derived from actual medical exam questions. In addition to correct answers, some of these datasets include detailed explanations. This enhances the model’s ability to select correct answers and improves its understanding and reasoning about complex medical information.
@@
Human Preference Data, such as Zhongjing_rlhf [179] and MedicalGPT [174], consist of a question, a chosen response, and a rejected response. The chosen response aligns with human values (more like a doctor’s response), whereas the rejected response is not consistent with human values compared to the chosen response. These datasets typically require manual labeling or verification, resulting in a smaller number of public datasets of this type compared to other types. Such datasets are used in the human alignment training stage of medical LLMs to ensure that the model’s responses align with human linguistic conventions and exhibit characteristics that are doctor-like, patient-friendly, harmless, and professional.
@@
Knowledge Graph represents concepts, entities, and their interrelationships in the real world in a structured format, aligning information more closely with human cognition. It enhances the organization, management, and comprehension of vast data [64, 41, 68, 157]. Medical knowledge graphs usually contain rich structured medical knowledge with low data noise. This type of data can motivate medical LLMs to more accurately process medical terminology and causal relationships, thus improving models’ performance in reasoning medical tasks. Common medical knowledge graphs include UMLS [96], CMeKG [17], and BIOS[13]. Since these are not in QA pairs or plain text format, a carefully designed data transformation pipeline is required to convert them.
@@
NLP Task datasets include Named Entity Recognition (NER) [87, 111], Relation Extraction (RE) [191, 125], Causal Relation Extraction (CRE) [178], Text Classification (TC) [108, 193], Semantic Textual Similarity (STS) [159, 183], Natural Language Inference (NLI) [4], and others. These datasets enhance model performance in various medical tasks and instruction following capabilities. Me-LLaMA [171] and Taiyi [102] utilize these conventional medical NLP datasets to improve their models’ generalization capabilities. It is important to note that these data generally cannot be used directly for training medical LLMs and must be transformed into plain text or question-answer pair format by data processing.
@@
2.1.2 Public Medical Corpus
@@
Publicly accessible online medical data is a primary source for training medical LLMs. The data encompasses diverse medical information from various regions, populations, and platforms. While it offers extensive coverage and a wide range of medical knowledge, it often presents challenges related to privacy and data quality, necessitating further cleaning and processing. Here, we summarize six common data sources for medical LLMs: medical encyclopedias, medical textbooks, medical academic literature, medical websites, medical guidelines, and package inserts.
@@
Medical Encyclopedias are comprehensive reference resources, typically available online, that systematically document a wide range of medical knowledge, including diseases, symptoms, diagnostic methods, treatments, and related medical technologies [48]. Compiled by professional medical groups, academic institutions, or authoritative publishers, their content undergoes rigorous review and regular updates to ensure accuracy and currency. The structured organization of these encyclopedias facilitates efficient data retrieval and processing. Notable medical encyclopedia websites include MedlinePlus [109], Mayo Clinic [107], and WebMD [161]. Medical LLMs often use encyclopedias as a training data source to enhance their medical knowledge [22, 156]. Although the format of encyclopedia content is suitable for the pretraining of these models, further data processing and transformation are necessary for instruction fine-tuning.
@@
Medical Textbooks provide a systematic introduction to medical knowledge and skills, serving as comprehensive references for medical students, physicians, nurses, and other healthcare professionals [140, 77]. They cover a wide range of topics from basic sciences (e.g., anatomy, physiology, biochemistry) to clinical practice (e.g., internal medicine, surgery, pediatrics, obstetrics and gynecology). Authored by experts and peer-reviewed, these textbooks ensure the accuracy and authority of medical information. Common sources include FreeBooks4Doctors [43], AccessMedicine [105], and Bookshelf [112]. Similar to medical encyclopedias, medical textbooks offer extensive, specialized, and structured information. Consequently, medical LLMs [89, 153, 189, 121] utilize these textbooks as training data to incorporate professional and precise medical knowledge. However, the specialized terminology and complex linguistic structures in medical textbooks can limit their accessibility for users without a medical background.
@@
Medical Academic Literature encompasses specialized publications that research, explore, and discuss various fields of medicine [83]. This literature includes peer-reviewed journal articles, scholarly books, conference papers, and research reports. Covering specialties such as clinical medicine, biomedical sciences, public health, pharmacology, and physiology, these publications provide up-to-date medical knowledge and evidence for researchers, clinicians, students, and allied professionals. Common sources include PubMed [110], Embase [39], and ScienceDirect [40]. Due to its specialization, accuracy, and currency, medical academic literature is frequently used to train medical LLMs [24, 171, 167], enhancing their medical expertise. However, the specialized terminology and technical content may make these models produce responses that are not easily understood by the general public. Additionally, acquiring academic literature data requires consideration of copyright issues.
@@
Medical Websites encompass various online platforms, such as forums, consultations, news, and educational resources. Specifically, medical consultation platforms allow patients to communicate with doctors online for medical advice, diagnoses, and treatment plans, enhancing the convenience and accessibility of healthcare services. These platforms also offer doctors a flexible way to provide care more efficiently. Common medical websites include 98point6 [1], Ding Xiang Yuan [38], Teladoc Health [139], and Haodf [56]. For training medical LLMs (LLMs), conversation data from medical consultation platforms is particularly valuable. This data is generally more accessible to the public compared to academic literature and textbooks due to its real-world, colloquial nature. Additionally, the multi-turn QA format helps improve the models’ ability to process contextual information. Several models, such as MMedLM [121], ChatDoctor [91], and GPT-Doctor [155], use this type of data to align their response styles with those of real doctors. However, data from medical consultation platforms must be collected with strict adherence to privacy regulations to ensure patient information security. The quality of this data can vary, with some information being inaccurate or incomplete. Therefore, it is essential to screen and clean the data to maintain the accuracy and reliability of the medical information used for training.
@@
Package Inserts are documents that provide comprehensive information about medications and their proper use [45]. Targeted at patients, doctors, and other healthcare professionals, it promotes the safe and effective use of medications. Package inserts typically include details such as the drug name and ingredients, usage and dosage, indications, contraindications, warnings and precautions, drug interactions, and side effects. Common sources for drug insert data include RxList [126] and Drugs [34]. These documents contain rich information that can enhance medical LLMs’ understanding of drugs and support decision-making processes. However, there are limitations. Package inserts offer standardized medication instructions and may not account for individual patient conditions or concurrent medication use, which can affect the ability of medical LLMs to provide personalized healthcare and treatment recommendations.
@@
2.1.3 Professional Medical Organization Corpus
@@
Data from specialized medical institutions, such as hospitals and healthcare centers, is a crucial information source for training medical LLMs. These data include patients’ medical histories, diagnostic information, treatment plans, and doctor-patient interactions. Here, we summarize four commonly used professional medical organization corpus sources: Clinical Notes, Electronic Health Records, Medical Prescriptions, and Doctor-Patient Conversations.
@@
Clinical Notes are records created by healthcare professionals during the consultation, diagnosis, treatment, and monitoring of patients. Typically prepared by doctors, nurses, and therapists, these notes include patient information, symptom descriptions, diagnoses, assessments, treatment plans, and progress notes [132]. MIMIC [75] is a publicly accessible database of clinical notes, offering diverse and comprehensive data covering the entire course of a patient’s medical care. These data will enhance medical LLMs’ abilities in clinical decision-making and case understanding. For instance, models like Zhongjing [179] and Me-LLaMA [171] use clinical notes for continued pretraining to enrich their knowledge and understand real-world clinical scenarios and reasoning. Meanwhile, Clinical Camel [144] aims to enable LLMs to synthesize plausible clinical records.
@@
Electronic Health Records (EHRs) are digital versions of patients’ medical records, offering a more comprehensive overview than clinical notes [63, 15]. EHRs include medical history, diagnoses, medication regimens, immunization dates, allergy information, radiology images, and laboratory test results. This real-time, patient-centric record system provides authorized users with immediate, secure access to patient information. Publicly accessible sources for EHRs include eICU [119], MIMIC-IV [74], and CPRD [61]. Given the close relation to medical practice and the multidimensional nature of the data, some works [179, 150] have utilized EHR data to train healthcare domain models, enhancing their usefulness and authenticity. However, different healthcare organizations may use varying EHR systems and record formats, necessitating additional processing to standardize the data. Furthermore, EHRs contain personal medical information that must be anonymized to protect patient privacy.
@@
Medical Prescription is an official directive from a physician or authorized healthcare professional to a pharmacist, detailing the medication to be dispensed to a patient, including the specific drug, dosage, frequency, and administration method [57]. These data reflect real-world prescription practices, providing valuable and authentic information to enhance models’ decision-making in healthcare. Some hospitals have integrated electronic medical prescriptions into their EHR systems, significantly facilitating scientific research on this data. Qibo [189] utilizes Traditional Chinese Medicine (TCM) prescriptions as pretraining corpus to incorporate TCM knowledge into medical LLMs. Medical Prescription data is relatively scarce and generally requires collaboration with hospitals to access. Additionally, these datasets may exhibit biases, with certain hospitals or physicians favoring specific drugs or treatments. Errors, omissions, or inaccuracies, such as misspelled drug names or incorrect dosages, can also compromise the effectiveness of model training and application.
@@
Doctor-Patient Conversation data encompasses real-life QA and interactions between doctors and patients, including symptom descriptions, diagnostic information, and patient feedback [104]. These conversations typically involve multiple rounds of communication, with doctors actively asking questions. Such data can enhance the capabilities of medical LLMs in assisting diagnosis, providing treatment suggestions, asking relevant questions, and understanding context. HuatuoGPT [188] uses real medical conversation data in serving instruction fine-tuning to enhance long conversations for medical LLMs. However, accessing this data is challenging due to privacy concerns from hospitals and patients, and the colloquial and emotional nature of these conversations requires future processing.
@@ 2.1.4 Synthetic Data
Synthetic data is generated by prompting an LLM to create content based on its internal knowledge or contextual information [134]. The common general synthetic datasets are Alpaca-52k [136], and ShareGPT [25], which are in QA format. These datasets aim to align customized LLMs with state-of-the-art LLMs at minimal cost. In the medical domain, several LLMs utilize state-of-the-art models to generate training datasets, constructing both single-turn and multi-turn QA formats. Related prompt templates are presented in the GitHub repository.
@@
For single-turn QA data, ChatMed-Consult [55] takes medical questions from the internet as input to ChatGPT and then concatenates its outputs to form an instruction fine-tuning dataset. HuatuoGPT [188] employs a self-instruct [160] method, generating a medical instruction dataset by manually creating seed instructions based on specific roles and use cases. Alpacare [195] initially creates a small set of clinician-curated seed tasks, considering four dimensions: topic, view, type, and difficulty level. Each task includes detailed information specific to these dimensions, instructions, and possibly corresponding inputs. GPT-4 then generates new instructions based on these seed tasks, filtering out highly similar ones. Finally, ChatGPT answers each generated instruction to compose the MedInstruct-52k instruction fine-tuning dataset. For multi-turn QA data, HuatuoGPT employs two ChatGPTs: one simulates a doctor and the other as a patient. Real data is utilized to stimulate conversations between these two ChatGPTs, resulting in a multi-turn QA dataset.
@@
The quality of the generated data depends on the effectiveness of the prompts. Well-designed prompts can significantly stimulate the intrinsic knowledge of LLMs, resulting in higher-quality data. However, since this data is not derived from real-world sources, manual checking and correction are typically required to enhance its quality.
2.2 Data Processing